An aeroplane heads due north at 500km/h. It experiences a 80km/h cross wind flowing in the direction N60 degrees E


find the true velocity of the aeroplane

determine the died of the aeroplane.


1
Expert's answer
2022-06-08T14:44:00-0400

a.


"\vec{v}=500\vec{j}""\vec{u}=80\sin60\degree \vec{i}+80\cos60\degree\vec{j}""\vec{v}_{res}=\vec{v}+\vec{u}""=40\sqrt{3}\vec{i}+540\vec{j}""|\vec{v}_{res}|=\sqrt{(40\sqrt{3})^2+(540)^2}=20\sqrt{741}(km/h)"



b.


"\tan \theta=\dfrac{540}{40\sqrt{3}}=4.5\sqrt{3}"

"\theta=\tan^{-1}(4.5\sqrt{3})\approx82.7\degree"

An aeroplane travels in the direction N "82.7\degree"E at "20\sqrt{741}" km/h.